# Enhancing Performance of Explainable AI Models with Constrained Concept Refinement

### v1

@@ C.4 Proof of Lemmas
C.4.1 Proof of Lemma 1
Lemma 1 directly follows from Lemma 2 and Lemma 3.
C.4.2 Proof of Lemma 2
@@
Note that ${\mathbf{z}}\sim\mathcal{N}(\mathbf{0},{\bm{I}}_{d\times d})$. This implies that $\Pi^{\perp}_{{\bm{D}}^{*}_{S}}{\mathbf{z}}$ is independent of $\Pi_{{\bm{D}}^{*}_{S}}{\mathbf{z}}$ [1], which in turn entails that $\langle\Pi^{\perp}_{{\bm{D}}^{*}_{S}}{\mathbf{x}},{\mathbf{z}}\rangle$ is independent of $\langle\Pi_{{\bm{D}}^{*}_{S}}{\mathbf{x}},{\mathbf{z}}\rangle$ and the events $\{\langle{\mathbf{d}}^{*}_{i},{\mathbf{z}}\rangle=r_{i}\quad\forall i\in S\}$. Therefore, we have:
@@
The equality $(a)$ is due to the observation that $\Pi^{\perp}_{{\bm{D}}^{*}_{S}}{\mathbf{z}}$ is indeed distributed as $\mathcal{N}(\mathbf{0},{\bm{I}}_{(d-k)\times(d-k)})$ after a simple dimensionality reduction [1]. The resulting random variable $\langle\Pi^{\perp}_{{\bm{D}}^{*}_{S}}{\mathbf{x}},\Pi^{\perp}_{{\bm{D}}^{*}_{S}}{\mathbf{z}}\rangle$ is distributed as $\mathcal{N}(0,\|\Pi^{\perp}_{{\bm{D}}^{*}_{S}}{\mathbf{x}}\|_{2}^{2})$, leading to the equality $(a)$. To sum up, we have
@@ C.4.3 Proof of Lemma 3
We will use induction to show that $S_{k}$ corresponds to the indices selected by IP-OMP after $k$ iterations. At the first iteration, according to [14], IP-OMP simply selects the index $i=\arg\max_{j\in[n]}|\langle{\mathbf{d}}_{j},{\mathbf{x}}\rangle|/\|{\mathbf{d}}_{j}\|_{2}\|{\mathbf{x}}\|_{2}=\arg\max_{j\in[n]}|\langle{\mathbf{d}}_{j},{\mathbf{x}}\rangle|$, where the second equality is due to $\|{\mathbf{d}}_{j}\|_{2}=1$ for all $j\in[n]$. As a result we have $S_{1}=\{i\}$ and the base case is established.
@@
Finally, when ${\bm{D}}={\bm{D}}^{*}$, one can verify that $\langle{\mathbf{d}}_{i},{\mathbf{x}}\rangle=\langle{\mathbf{d}}^{*}_{i},{\bm{D}}^{*}\mathbf{\beta}\rangle=0$ for $i\not\in S^{*}$. This implies that $S_{k}=S^{*}$. $\hfill\qed$
C.4.4 Proof of Lemma 4
@@
C.4.5 Proof of Lemma 5
The proof strategy we adopt here is similar with the proof of Lemma 3.1 in [34]. However, we consider a different generative model in this paper, so we present the full proof for the purpose of completeness. We will show that for $i\in S^{*}$, $|\langle{\mathbf{d}}_{i},{\mathbf{x}}\rangle|>\gamma/2$ and for $i\not\in S^{*}$, $|\langle{\mathbf{d}}_{i},{\mathbf{x}}\rangle|<\gamma/2$, which will immediately result in $S=\operatorname*{arg\,max}_{T\subseteq[n],|T|\leq k}\sum_{i\in T}|\langle{\mathbf{d}}_{i},{\mathbf{x}}\rangle|=S^{*}$, which proves Lemma 5.
@@ C.4.6 Proof of Lemma 6
C.4.7 Proof of Lemma 7
The proof of Lemma 7 consists of two steps:
@@ C.4.7 Proof of Lemma 7
Combining these two steps and taking the union bound will complete the proof of Lemma 7.
@@
$\sigma_{d}\left(\sum_{h\in Q_{i}}{\mathbf{x}}^{h}{\mathbf{x}}^{h\top}\right)$ $=\sigma_{d}\left(\sum_{h\in Q_{i}}{\bm{D}}^{*}\mathbf{\beta}^{h}\mathbf{\beta}^{h\top}{\bm{D}}^{*\top}\right)$
$=\sigma_{d}\left(\sum_{h\in Q_{i}}\mathbf{\beta}^{h}\mathbf{\beta}^{h\top}\right)$
The last equation is due to the fact that ${\bm{D}}^{*}$ is a full-rank orthogonal matrix. Consider the vector $\mathbf{\beta}^{h}$ for $h\in Q_{i}$. According to our generative model and the definition of $Q_{i}$, we have $\mathbf{\beta}^{h}_{i}\neq 0$, and its support outside $i$ is selected uniformly over all $(k-1)$-element subsets of $[n]\backslash\{i\}$. Each nonzero entry of $\mathbf{\beta}^{h}$ has a zero mean, variance of $\sigma^{2}$, and an absolute value bounded between $\gamma$ and $\Gamma$. Therefore, its covariance matrix defined as ${\bm{\Sigma}}:=\mathbb{E}\left[\mathbf{\beta}^{h}{\mathbf{\beta}^{h}}^{\top}\right]$ is diagonal, with its $(i,i)$-th entry corresponding to $\sigma^{2}$, and its $(j,j)$-th entry (with $j\not=i$) corresponding to $\frac{(k-1)\sigma^{2}}{n}$. We claim that when $|Q_{i}|$ is sufficiently large, $\sum_{h\in Q_{i}}{\mathbf{x}}^{h}{\mathbf{x}}^{h\top}$ will concentrate around $|Q_{i}|{\bm{\Sigma}}$. Specifically, we will use the following well-established result:
Theorem 5 ([60]).
@@
To invoke Theorem 5, we note that $\mathbf{\beta}^{h}$ is indeed zero-mean and sub-Gaussian. Therefore, upon setting ${\mathbf{w}}=\mathbf{\beta}^{h}$, we notice that for any unit-norm ${\mathbf{q}}$:
$\|\langle\mathbf{\beta}^{h},{\mathbf{q}}\rangle\|^{2}_{\psi_{2}}\leq\sum_{j=1}^{d}{\mathbf{q}}_{j}^{2}\|\mathbf{\beta}^{h}_{j}\|^{2}_{\psi_{2}}\leq C_{0}\sigma^{2},$
@@
$\mathbb{E}\left[\left\langle\mathbf{\beta}^{h},{\mathbf{q}}\right\rangle^{2}\right]={\mathbf{q}}^{\top}\Sigma{\mathbf{q}}\geq\frac{(k-1)\sigma^{2}}{n}.$
By setting $K^{2}=\frac{C_{0}n}{k-1}$, we can invoke Theorem 5 with $u=n$ and conclude that, with probability at least $1-2\exp\{-n\}$:
$\sigma_{d}\left(\sum_{h\in Q_{i}}\mathbf{\beta}^{h}\mathbf{\beta}^{h\top}\right)$ $\geq|Q_{i}|\sigma_{d}\left({\bm{\Sigma}}\right)-\frac{Cn\sigma}{k-1}\sqrt{2n|Q_{i}|}=\frac{|Q_{i}|(k-1)\sigma^{2}}{n}-\frac{Cn\sigma}{k-1}\sqrt{2n|Q_{i}|},$ (23)
$\sigma_{1}\left(\sum_{h\in Q_{i}}\mathbf{\beta}^{h}\mathbf{\beta}^{h\top}\right)$ $\leq|Q_{i}|\sigma_{1}\left({\bm{\Sigma}}\right)+\frac{Cn\sigma}{k-1}\sqrt{2n|Q_{i}|}=|Q_{i}|\sigma^{2}+\frac{Cn\sigma}{k-1}\sqrt{2n|Q_{i}|},$ (24)
@@
We can then invoke Theorem 6 with $\delta=\frac{1}{2}$ to obtain
@@
$\frac{1}{m}\sigma_{d}\left(\sum_{h\in Q_{i}}\mathbf{\beta}^{h}\mathbf{\beta}^{h\top}\right)$ $\geq\frac{|Q_{i}|(k-1)\sigma^{2}}{nm}-\frac{Cn\sigma}{(k-1)m}\sqrt{2n|Q_{i}|}$
@@
$\frac{1}{m}\sigma_{d}\left(\sum_{h\in Q_{i}}\mathbf{\beta}^{h}\mathbf{\beta}^{h\top}\right)\geq\frac{k(k-1)\sigma^{2}}{2n^{2}}-\frac{k(k-1)\sigma^{2}}{4n^{2}}=\frac{k(k-1)\sigma^{2}}{4n^{2}}.$
@@
$\frac{1}{m}\sigma_{1}\left(\sum_{h\in Q_{i}}\mathbf{\beta}^{h}\mathbf{\beta}^{h\top}\right)\leq|Q_{i}|\sigma^{2}+\frac{Cn\sigma}{k-1}\sqrt{2n|Q_{i}|}\leq\frac{4k\sigma^{2}}{n}.$
@@ D.1 Experiments on Generative Model
We generate samples according to the described generative model in 3, we construct the input data ${\mathbf{x}}$ or $\{{\mathbf{x}}^{h}\}_{h=1}^{m}$ by sampling each non-zero entry from a uniform distribution over the interval $[\gamma,\Gamma]$. The matrix ${\bm{D}}^{*}$ is chosen to be a randomly generated orthonormal matrix, and we construct ${\bm{D}}_{\mathrm{init}}$ by setting ${\bm{D}}_{\mathrm{init}}={\bm{D}}^{*}+{\bm{E}}$, where each column of ${\bm{E}}$ is uniformly drawn from the set $\{{\mathbf{e}}\mid\|{\mathbf{e}}\|_{2}\leq\rho\}$. For the results shown in Figure 4, we set $d=10$, $k=5$, $\rho=0.2$, $\gamma=0.5$, and $\Gamma=1$. The first column of Figure 4 illustrates the scenario corresponding to Theorem 3, in which only a single input feature ${\mathbf{x}}$ is available. Here, we choose $n=8$ and $\eta=10^{-2}$. The plot showing $\mathcal{L}(\tilde{f}_{{\bm{D}}^{(t)}})$ confirms the linear convergence of the loss function to zero, in agreement with Theorem 3. However, as previously noted, this setting does not ensure exact recovery for each queried feature. Indeed, as shown in the right panel of the first row, $\|{\bm{D}}^{(t)}-{\bm{D}}^{*}\|_{1,2}$ does not converge to zero. Even if we only consider columns that are activated at each iteration, $\|{\bm{D}}^{(t)}_{S}-{\bm{D}}^{*}_{S}\|_{1,2}$ also remains nonzero, indicating that optimizing $\mathcal{L}$ with a single ${\mathbf{x}}$ is insufficient to recover the ground-truth dictionary ${\bm{D}}^{*}$.
In the second column of Figure 4, we apply projected gradient descent to minimize $\mathcal{L}_{m}$, as defined in Equation 5, under the assumption that ${\bm{D}}^{*}$ is rank-deficient. Specifically, we set $n=8<d$ and $\eta=10^{-1}$. Evidently, $\|{\bm{D}}^{(t)}-{\bm{D}}^{*}\|_{1,2}$ does not converge to zero because the component of ${\bm{D}}^{(t)}-{\bm{D}}^{*}$ orthogonal to the column space of ${\bm{D}}^{*}$ remains unaffected throughout the iterations. In contrast, the third column of Figure 4 corresponds to the setting $n=d$, which guarantees the full-rankness of ${\bm{D}}^{*}$ is full rank. In this case, $\|{\bm{D}}^{(t)}_{S}-{\bm{D}}^{*}_{S}\|_{1,2}$ converges to zero, thus supporting the conclusion of Theorem 4.
@@ D.2 More Details on Algorithm 2
D.3 More Experiments on Interpretability
In this section, we expand on the discussion of CCR’s interpretability from Section 4.2 by conducting a comprehensive case study across all five datasets used to evaluate CCR. For each dataset, we select three representative image-label pairs and present the ten highest-ranking concept scores on the left, alongside their corresponding weights associated with the predicted label on the right. Based on our case study, we derive several key observations:

### v2

@@ C.4 Proof of Lemmas
C.4.1 Proof of Lemma 3.1
Lemma 3.1 directly follows from Lemma C.1 and Lemma C.2.
C.4.2 Proof of Lemma C.1
@@
Note that ${\mathbf{z}}\sim\mathcal{N}(\mathbf{0},{\bm{I}}_{d\times d})$. This implies that $\Pi^{\perp}_{{\bm{D}}^{*}_{S}}{\mathbf{z}}$ is independent of $\Pi_{{\bm{D}}^{*}_{S}}{\mathbf{z}}$ (Anderson et al., 1958), which in turn entails that $\langle\Pi^{\perp}_{{\bm{D}}^{*}_{S}}{\mathbf{x}},{\mathbf{z}}\rangle$ is independent of $\langle\Pi_{{\bm{D}}^{*}_{S}}{\mathbf{x}},{\mathbf{z}}\rangle$ and the events $\{\langle{\mathbf{d}}^{*}_{i},{\mathbf{z}}\rangle=r_{i}\quad\forall i\in S\}$. Therefore, we have:
@@
The equality $(a)$ is due to the observation that $\Pi^{\perp}_{{\bm{D}}^{*}_{S}}{\mathbf{z}}$ is indeed distributed as $\mathcal{N}(\mathbf{0},{\bm{I}}_{(d-k)\times(d-k)})$ after a simple dimensionality reduction (Anderson et al., 1958). The resulting random variable $\langle\Pi^{\perp}_{{\bm{D}}^{*}_{S}}{\mathbf{x}},\Pi^{\perp}_{{\bm{D}}^{*}_{S}}{\mathbf{z}}\rangle$ is distributed as $\mathcal{N}(0,\|\Pi^{\perp}_{{\bm{D}}^{*}_{S}}{\mathbf{x}}\|_{2}^{2})$, leading to the equality $(a)$. To sum up, we have
@@
We will use induction to show that $S_{k}$ corresponds to the indices selected by IP-OMP after $k$ iterations. At the first iteration, according to (Chattopadhyay et al., 2024), IP-OMP simply selects the index $i=\arg\max_{j\in[n]}|\langle{\mathbf{d}}_{j},{\mathbf{x}}\rangle|/\|{\mathbf{d}}_{j}\|_{2}\|{\mathbf{x}}\|_{2}=\arg\max_{j\in[n]}|\langle{\mathbf{d}}_{j},{\mathbf{x}}\rangle|$, where the second equality is due to $\|{\mathbf{d}}_{j}\|_{2}=1$ for all $j\in[n]$. As a result we have $S_{1}=\{i\}$ and the base case is established.
@@
Finally, when ${\bm{D}}={\bm{D}}^{*}$, one can verify that $\langle{\mathbf{d}}_{i},{\mathbf{x}}\rangle=\langle{\mathbf{d}}^{*}_{i},{\bm{D}}^{*}\bm{\beta}\rangle=0$ for $i\not\in S^{*}$. This implies that $S_{k}=S^{*}$. $\hfill\qed$
C.4.4 Proof of Lemma C.3
@@
C.4.5 Proof of Lemma C.4
The proof strategy we adopt here is similar with the proof of Lemma 3.1 in (Liang et al., 2022). However, we consider a different generative model in this paper, so we present the full proof for the purpose of completeness. We will show that for $i\in S^{*}$, $|\langle{\mathbf{d}}_{i},{\mathbf{x}}\rangle|>\gamma/2$ and for $i\not\in S^{*}$, $|\langle{\mathbf{d}}_{i},{\mathbf{x}}\rangle|<\gamma/2$, which will immediately result in $S=\operatorname*{arg\,max}_{T\subseteq[n],|T|\leq k}\sum_{i\in T}|\langle{\mathbf{d}}_{i},{\mathbf{x}}\rangle|=S^{*}$, which proves Lemma C.4.
@@
C.4.7 Proof of Lemma C.6
The proof of Lemma C.6 consists of two steps:
@@
Combining these two steps and taking the union bound will complete the proof of Lemma C.6.
@@
$\sigma_{d}\left(\sum_{h\in Q_{i}}{\mathbf{x}}^{h}{\mathbf{x}}^{h\top}\right)$ $=\sigma_{d}\left(\sum_{h\in Q_{i}}{\bm{D}}^{*}\bm{\beta}^{h}\bm{\beta}^{h\top}{\bm{D}}^{*\top}\right)$
$=\sigma_{d}\left(\sum_{h\in Q_{i}}\bm{\beta}^{h}\bm{\beta}^{h\top}\right)$
The last equation is due to the fact that ${\bm{D}}^{*}$ is a full-rank orthogonal matrix. Consider the vector $\bm{\beta}^{h}$ for $h\in Q_{i}$. According to our generative model and the definition of $Q_{i}$, we have $\bm{\beta}^{h}_{i}\neq 0$, and its support outside $i$ is selected uniformly over all $(k-1)$-element subsets of $[n]\backslash\{i\}$. Each nonzero entry of $\bm{\beta}^{h}$ has a zero mean, variance of $\sigma^{2}$, and an absolute value bounded between $\gamma$ and $\Gamma$. Therefore, its covariance matrix defined as ${\bm{\Sigma}}:=\mathbb{E}\left[\bm{\beta}^{h}{\bm{\beta}^{h}}^{\top}\right]$ is diagonal, with its $(i,i)$-th entry corresponding to $\sigma^{2}$, and its $(j,j)$-th entry (with $j\not=i$) corresponding to $\frac{(k-1)\sigma^{2}}{n}$. We claim that when $|Q_{i}|$ is sufficiently large, $\sum_{h\in Q_{i}}{\mathbf{x}}^{h}{\mathbf{x}}^{h\top}$ will concentrate around $|Q_{i}|{\bm{\Sigma}}$. Specifically, we will use the following well-established result:
Theorem C.7 ((Vershynin, 2018)).
@@
To invoke Theorem C.7, we note that $\bm{\beta}^{h}$ is indeed zero-mean and sub-Gaussian. Therefore, upon setting ${\mathbf{w}}=\bm{\beta}^{h}$, we notice that for any unit-norm ${\mathbf{q}}$:
$\|\langle\bm{\beta}^{h},{\mathbf{q}}\rangle\|^{2}_{\psi_{2}}\leq\sum_{j=1}^{d}{\mathbf{q}}_{j}^{2}\|\bm{\beta}^{h}_{j}\|^{2}_{\psi_{2}}\leq C_{0}\sigma^{2},$
@@
$\mathbb{E}\left[\left\langle\bm{\beta}^{h},{\mathbf{q}}\right\rangle^{2}\right]={\mathbf{q}}^{\top}\Sigma{\mathbf{q}}\geq\frac{(k-1)\sigma^{2}}{n}.$
By setting $K^{2}=\frac{C_{0}n}{k-1}$, we can invoke Theorem C.7 with $u=n$ and conclude that, with probability at least $1-2\exp\{-n\}$:
$\sigma_{d}\left(\sum_{h\in Q_{i}}\bm{\beta}^{h}\bm{\beta}^{h\top}\right)$ $\geq|Q_{i}|\sigma_{d}\left({\bm{\Sigma}}\right)-\frac{Cn\sigma}{k-1}\sqrt{2n|Q_{i}|}=\frac{|Q_{i}|(k-1)\sigma^{2}}{n}-\frac{Cn\sigma}{k-1}\sqrt{2n|Q_{i}|},$ (24)
$\sigma_{1}\left(\sum_{h\in Q_{i}}\bm{\beta}^{h}\bm{\beta}^{h\top}\right)$ $\leq|Q_{i}|\sigma_{1}\left({\bm{\Sigma}}\right)+\frac{Cn\sigma}{k-1}\sqrt{2n|Q_{i}|}=|Q_{i}|\sigma^{2}+\frac{Cn\sigma}{k-1}\sqrt{2n|Q_{i}|},$ (25)
@@
We can then invoke Theorem C.8 with $\delta=\frac{1}{2}$ to obtain
@@
$\frac{1}{m}\sigma_{d}\left(\sum_{h\in Q_{i}}\bm{\beta}^{h}\bm{\beta}^{h\top}\right)$ $\geq\frac{|Q_{i}|(k-1)\sigma^{2}}{nm}-\frac{Cn\sigma}{(k-1)m}\sqrt{2n|Q_{i}|}$
@@
$\frac{1}{m}\sigma_{d}\left(\sum_{h\in Q_{i}}\bm{\beta}^{h}\bm{\beta}^{h\top}\right)\geq\frac{k(k-1)\sigma^{2}}{2n^{2}}-\frac{k(k-1)\sigma^{2}}{4n^{2}}=\frac{k(k-1)\sigma^{2}}{4n^{2}}.$
@@
$\frac{1}{m}\sigma_{1}\left(\sum_{h\in Q_{i}}\bm{\beta}^{h}\bm{\beta}^{h\top}\right)\leq|Q_{i}|\sigma^{2}+\frac{Cn\sigma}{k-1}\sqrt{2n|Q_{i}|}\leq\frac{4k\sigma^{2}}{n}.$
@@ D.1 Experiments on Generative Model
We generate samples according to the described generative model in 3.2, we construct the input data ${\mathbf{x}}$ or $\{{\mathbf{x}}^{h}\}_{h=1}^{m}$ by sampling each non-zero entry from a uniform distribution over the interval $[\gamma,\Gamma]$. The matrix ${\bm{D}}^{*}$ is chosen to be a randomly generated orthonormal matrix, and we construct ${\bm{D}}_{\mathrm{init}}$ by setting ${\bm{D}}_{\mathrm{init}}={\bm{D}}^{*}+{\bm{E}}$, where each column of ${\bm{E}}$ is uniformly drawn from the set $\{{\mathbf{e}}\mid\|{\mathbf{e}}\|_{2}\leq\rho\}$. For the results shown in Figure 4, we set $d=10$, $k=5$, $\rho=0.2$, $\gamma=0.5$, and $\Gamma=1$. The first column of Figure 4 illustrates the scenario corresponding to Theorem 3.3, in which only a single input feature ${\mathbf{x}}$ is available. Here, we choose $n=8$ and $\eta=10^{-2}$. The plot showing $\mathcal{L}(\tilde{f}_{{\bm{D}}^{(t)}})$ confirms the linear convergence of the loss function to zero, in agreement with Theorem 3.3. However, as previously noted, this setting does not ensure exact recovery for each queried feature. Indeed, as shown in the right panel of the first row, $\|{\bm{D}}^{(t)}-{\bm{D}}^{*}\|_{1,2}$ does not converge to zero. Even if we only consider columns that are activated at each iteration, $\|{\bm{D}}^{(t)}_{S}-{\bm{D}}^{*}_{S}\|_{1,2}$ also remains nonzero, indicating that optimizing $\mathcal{L}$ with a single ${\mathbf{x}}$ is insufficient to recover the ground-truth dictionary ${\bm{D}}^{*}$.
In the second column of Figure 4, we apply projected gradient descent to minimize $\mathcal{L}_{m}$, as defined in Equation 6, under the assumption that ${\bm{D}}^{*}$ is rank-deficient. Specifically, we set $n=8<d$ and $\eta=10^{-1}$. Evidently, $\|{\bm{D}}^{(t)}-{\bm{D}}^{*}\|_{1,2}$ does not converge to zero because the component of ${\bm{D}}^{(t)}-{\bm{D}}^{*}$ orthogonal to the column space of ${\bm{D}}^{*}$ remains unaffected throughout the iterations. In contrast, the third column of Figure 4 corresponds to the setting $n=d$, which guarantees the full-rankness of ${\bm{D}}^{*}$ is full rank. In this case, $\|{\bm{D}}^{(t)}_{S}-{\bm{D}}^{*}_{S}\|_{1,2}$ converges to zero, thus supporting the conclusion of Theorem 3.4.
@@ D.2 More Details on Algorithm 2
The effectiveness of Algorithm 3 is shown in Figure 5, where we compare the correlation between concept embeddings before and after this process across various datasets.
@@
Figure 5: We calculate the correlation matrix ($\mathbf{D}^{\top}\mathbf{D}$) for dictionaries before and after Algorithm 3, and present them in the format of heatmaps. As can be seen, the proposed dispersion process effectively reduces the correlation between concept embeddings generated by CLIP.
D.3 Ablation Study
Figure 6: Ablation study for Algorithm 3 and Algorithm 4.
In this section, we present empirical findings from our ablation study. As a preliminary note, we emphasize that the comparison between CCR and its baseline shown in Figure 2 serves as a key ablation analysis for the CCR module as a whole. Here, we concentrate on evaluating the contributions of two critical components introduced in the preceding section: Atom Dispersion and Atom Projection. To this end, we individually exclude these two steps—Algorithm 3 and Algorithm 4—and compare the resulting models to the full CCR framework in terms of test accuracy, average sparsity, average column deviation, and maximum column deviation over the course of training.
As illustrated in Figure 6, the omission of either step leads to an improvement in performance. However, these gains incur different trade-offs. Specifically, removing Algorithm 3 results in a substantial increase in the average sparsity of the learned sparse codes. In contrast, eliminating Algorithm 4 causes the concept atoms to drift significantly from their initial CLIP embeddings, thereby compromising the interpretability of the model.
D.4 Hyperparameter Tuning
This section presents an empirical evaluation of two critical hyperparameters—namely, the hard-threshold parameter $\lambda$ and the radius bound $\rho$—which significantly influence the performance of Algorithm 2. The corresponding results are illustrated in Figure 7 and Figure 8.
As depicted in Figure 7, increasing the value of $\lambda$ results in sparser explanations by reducing the number of activated concepts. However, this sparsity comes at the cost of reduced test accuracy. Therefore, selecting an appropriate threshold entails balancing the trade-off between prediction accuracy and the level of explanation sparsity—a consideration of particular relevance in explainable AI applications.
It is important to note that including more concepts does not inherently ensure improved accuracy. This point is exemplified in the left column of Figure 8, which shows that expanding the search radius enables the CCR algorithm to identify solutions that are simultaneously sparser and more accurate. Furthermore, the empirical results suggest that the performance gains of CCR saturate beyond $\rho>0.1$, indicating that $\rho=0.1$ serves as a suitable choice for this hyperparameter.
Figure 7: Average sparsity and test accuracy for varying thresholds $\lambda$.
Figure 8: Average sparsity and test accuracy for varying radius bounds $\rho$.
D.5 More Experiments on Interpretability
In this section, we expand on the discussion of CCR’s interpretability from Section 4.2 by conducting a comprehensive case study across all five datasets used to evaluate CCR. For each dataset, we select three representative image-label pairs and present the ten highest-ranking concept scores on the left, alongside their corresponding weights associated with the predicted label on the right. To facilitate a comparison of explainability with and without CCR, we also report the corresponding results for the baseline model, which is obtained by setting $\eta_{\mathbf{D}}=0$ in Algorithm 2, presented at the end of this section. Based on our case study, we derive several key observations: